OpenStudy (anonymous):

squarrt6+2squarert6 : simplify the expression

OpenStudy (anonymous):

3sqrt6

OpenStudy (anonymous):

When adding and subtracting square roots with the same radicals, you keep that and just add the terms. \[\sqrt{6}+2\sqrt{6}\] \[3\sqrt{6}\]

OpenStudy (lgbasallote):

\[\sqrt 6 + 2\sqrt6\] simply think of this as something like \[x + 2x\] what is the value of that?

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

3sqrt6